OpenStudy (anonymous):

Solve equation please help 2-3e^(.4x)=-7

jhonyy9 (jhonyy9):

so like first step add to both sides (-2) and will get ...? can you calcule it ?

OpenStudy (anonymous):

-3e^(.4x)=-9

jhonyy9 (jhonyy9):

ok what you can doing now ?

OpenStudy (anonymous):

Divide by -3 e^(.4x)=3

jhonyy9 (jhonyy9):

ok and now ? to solve it for x you need us logarithm like ,,ln" because ,,ln" has the base of ,,e"

jhonyy9 (jhonyy9):

so than will be ...?

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (anonymous):

Am I supposed to use change of base property?

jhonyy9 (jhonyy9):

why ? so this will be : ln e^(4x) =ln 3

jhonyy9 (jhonyy9):

continue it please

OpenStudy (anonymous):

I don't know what to do next.

jhonyy9 (jhonyy9):

4xln e =ln 3 4x= (ln 3 / ln e) so but you know that the natural logarithm has a base of e abd you know again that log(a) a =1 than this mean that ln e =1 ok ?

OpenStudy (anonymous):

Ok

jhonyy9 (jhonyy9):

than will get 4x=ln 3 x=(ln 3)/4 right ?

jhonyy9 (jhonyy9):

do you understand it sure ?

OpenStudy (anonymous):

Yes sir thank you.

jhonyy9 (jhonyy9):

ok good luck bye was my pleasure

OpenStudy (anonymous):

@precal

jhonyy9 (jhonyy9):

so change of base for ??? why ?

OpenStudy (precal):